OpenStudy (anonymous):

Evaluate the expression when a = 2 and b = 4. 6a – 2b + 5

OpenStudy (anonymous):

9

OpenStudy (johnweldon1993):

So you would just plug in 2 for 'a' and 4 for 'b' 6a - 2b + 5 would become 6(2) - 2(4)+ 5 12 - 8 + 5 = ?

OpenStudy (arabpride):

6a – 2b + 5 6(2) - 2(4) = 4 multiply: 12 - 8 = 4 that makes sense right *^.^*

OpenStudy (anonymous):

would equal 9

OpenStudy (johnweldon1993):

That is correct @stephaniej1 :)

OpenStudy (anonymous):

yup thanks guys x))

OpenStudy (arabpride):

NP